How do u solve this question?
A study monitored the life spans of 129 robins over 4 years. The equation was developed to calculate the number of years y for x percent of the robin population to die.

y= 2-log(100-x)/.42
How many robins had died after 6 months?

Answer by stanbon(48535)

A study monitored the life spans of 129 robins over 4 years. The equation was developed to calculate the number of years y for x percent of the robin population to die.
y= 2-log(100-x)/.42
How many robins had died after 6 months?
Let y=(1/2) yr and solve for x:
0.5=2-log(100-x)/0.42
-1.5=-log(....)/0.42
log(100-x)=0.63
100-x=10^0.63
100-x=4.2657
x=95.73%
-------------
# of robins that died is 0.9573*129=123+
============
Cheers,
Stan H.